The girl has a speed of 7 m/s at the bottom of the ramp. Calculate the kinetic energy of the girl at the bottom of the ramp use the equation kinetic energy = 0.5 x mass x (speed)2

1
Expert's answer
2021-10-07T08:49:29-0400

According to the formula, the kinetic energy is:


"K = \dfrac{m\cdot (7m/s)^2}{2} = 24.5m"

Thus, the kinetic energy is "24.5m(J)" if mass in substituted in "kg" into this formula.